Find the midpoint and the distance between the points (6,-8) and (14,7).
Thanks for your help!!!

Answer by venugopalramana(3286)

Find the midpoint and the distance between the points (6,-8) and (14,7).
DISTANCE BETWEEN (X1,Y1) AND (X2,Y2) IS GIVEN BY
SQRT.{(X2-X1)^2+(Y2-Y1)^2}
AND THEIR MID POINT IS GIVEN BY {(X1+X2)/2,(Y1+Y2)/2}
SO WE GET DISTANCE =SQRT.{(14-6)^2+(7+8)^2}=SQRT(64+225)=SQRT.(289)
MIDPOINT IS {(6+14)/2,(-8+7)/2}=(10,-1/2)